# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v1

@@ IV-B2 Attention Mechanism
The discussion of attention in the current approach focuses on decoder-only LLM-based embedder. Casual Attention [148] is a common operation for decoder-based LLMs, which ensures that the language modeling can only refer to the prefix to predict the next token. However, extensive empirical studies [149, 117, 128] have shown that casual attention will degrade downstream task performance. Current approaches either (1) retain the casual attention and use other tricks (e.g., special pooling methods) or (2) convert it to bi-directional attention and make the model adapt to it.
Maintance the Casual Attention
@@
Therefore, if the model maintains casual attention, it is often necessary to resort to other tricks to mitigate the resulting negative effects. Representatives of these methods include SGPT [119], PromptEOL [114] and its variants [115, 116], and Echo [128]; please refer to Section IV-B1 for details.
@@
Considering the use of a mask matrix implementation in Transformer  [^4] https://github.com/huggingface/transformers, the most popular framework to instantiate LLMs, it is convenient and quick to remove the mask matrix during incremental fine-tuning. BeLLM [117] pioneered an attempt to change the casual attention of the last few layers in the decoder-only LLM to bi-directional attention. This trial is motivated by the fact that as the number of layers increases, the performance of downstream tasks does not improve monotonically, but there is a turning point. When fully converting the casual attention in decoder-only LLM to bi-directional, sufficient data or additional training tasks are required to adapt the LLM to the new attentional mechanism. The experimental results in [117] show that when training on small-scale data, changing the attention of only the last layer from casual to bi-directional can effectively improve the performance on the SemEval Benchmark; however, changing all the attentional layers directly to bi-directional will result in a significant decrease in performance. Further, [136] shows that LLMs are able to adaptively update the parameters and complete the conversion from casual attention to bi-attention without additional operations when using sufficient data (hundreds of datasets) for incremental fine-tuning. To adapt the LLM to use bi-direction attention without massive training data, [130] propose a self-supervised task, masked next token prediction (MNTP). MNTP combines the ideas of mask language modeling and next token prediction, where the token is first randomly masked off from the text, and the hidden state from the previous position of the mask token is used to predict the mask token.
Remark 1
When the input contains instructions or the LLMs are expected to accomplish both generation and embedding tasks, we refer the reader to the detailed ablation experiments in [134] to learn more.
IV-B3 Additional Projector
Adding projectors is the usual strategy in text embedding, and we divide the existing work according to the purpose of the projectors.
@@ IV-B3 Additional Projector
The hidden layer dimension (width) usually increases with the number of total parameters. The hidden states of T5-11B have 1024 dimensions, while the number is 4096 for the mainstream 7B decoder-only LLMs, which are both higher than the 768 dimensions for a bunch of BERT-based models. In practice, it is obvious that high-dimensional embedding dramatically increases the resource overhead for storage and inference, and a simple approach is to add a projector $g:\mathbb{R}^{d}\to\mathbb{R}^{m},m<d$ after pooling to obtain the low-dimensional embedding. Sentence-T5 [113] and GTR [118] empirically demonstrate that the performance on STS, classification, and retrieval tasks can be improved by scaling the embedder’s parameter only without increasing the text embedding dimension. A linear layer without bias is used to transform the hidden state dimensions of different models to the same 768-dimensional text embedding. GirtLM [134] also uses a linear layer to map the 4096-dimensional original embedding to 1024 and finds that the performance of MTEB would be slightly degraded.
@@
Similar to the purpose of dimensionality reduction, converting text embedding to sparse representations can improve the inference efficiency of the model [150] and enhance its performance on downstream tasks such as long document retrieval [58]. For BERT-based embedders, text embeddings can be mapped to the vocabulary-length logits through the projector for mask language modeling; then, the logits can be regarded as the text representation. To sparsify the representation, common methods include gate mechanism [151], top-k masking [152], and regularization terms [153]. Similarly, text embeddings from LLM-based embedders can go through the decoder layer and use a similar idea for sparsification. Specifically, PromptReps [124] uses the prompt similar to PromptEOL to obtain text embeddings and introduce ReLU function, log-saturation [150] and top-k masking to obtain the sparse representation for documents. Mistral-SPLADE [126] use the prompt same to Echo [128] to obtain text embeddings and follow the SPLADE recipe [150] to fine-tune a Mistral-based embedder. [154] finds that the embedding output from LLM-based embedders, which is fine-tuned by contrastive learning, still produces high-quality sparse representations via the decoding layer and simple top-k masking.
@@
InBedder [133] proposes using question-answer pairs with short-length answers to train text embedder. Specifically, the context and question are filled into a prompt template, and the LLM is fine-tuned to predict the short-length answer with the standard next-token prediction. In this way, the hidden state used to predict the first token is shown to work as high-quality text embedding, while the fine-tuned embedder is shown to have a good instruction-following ability. Thus, this is a special case where the decoder layer is reserved for projectors to be used in training text embedders. Search-Adaptor [155] proposes the use of additional projectors for customizing commercial text embedding interfaces.
IV-B4 Parameter-Efficient Fune-Tuning Module
The current parameter scales used for treating LLM-based Embedder are around 7B, e.g., Mistral-7B and LLaMA3-8B are a few of the commonly used components. Interestingly, the only few attempts to use LLM with larger LLMs lead to even worse performance [134]. Under the popular half-precision (TF16 or BF16) setting, one 7B LLM can occupy $\sim$60GB RAM for full-parameter fine-tuning or $\sim$16G RAM for parameter-efficient fine-tuning using LoRA, which both can be accomplished on 2 V100 or A100 GPUs. Thus, whether or not parameter-efficient fine-tuning techniques are used depends primarily on the amount of data used, not on resource constraints. Some works introduce BitFit [156], or LoRA [157] as the PEFT modules and fine-tune using single datasets, such as Wiki1M [130], SNLI [119], or MSMARCO [120], in the belief that a small-size data can stimulate the semantic generalization capabilities that LLMs themselves possess. The other works use full parameter tuning and (multiple-stage) fine-tuning based on hundreds of datasets, which was sufficient to allow large changes to the model parameters and avoid overfitting.
IV-C Optimization
@@
IV-C1 Prompt Engineering-Based Methods (PE)
Prompt engineering leverages the inherent capabilities of large language models without additional fine-tuning. By carefully crafting input prompts, practitioners can guide the model to produce desirable embeddings tailored to specific tasks. Several notable methods have been developed. PromptEOL [114] utilizes in-context learning to enhance sentence embedding performance. Echo [128] populates individual texts into a prompt template twice and obtains the textual embedding from the second occurrence position of the text. The PromptReps [124] method extends the use of prompt engineering to generate retrieval-specific embeddings. By inputting documents and queries along with relevant prompts into the LLM, it utilizes the embedding at the <eos> position as a feature for retrieval tasks. MetaEOL [115] introduces an unsupervised text embedding technique known as Meta-Task Prompting with Explicit One-Word Limitation. In the study on PromptSTH/SUM [116], the authors introduce two novel prompt engineering techniques—Pretended Chain of Thought and Knowledge Enhancement—which enhance the expressiveness of LLM embeddings and are effective across diverse LLMs. BeLLM [117] explores the impact of backward dependencies in LLMs and then proposes a backward dependency enhanced large language model (BeLLM) that transforms certain attention layers from unidirectional to bidirectional. This modification allows the model to learn more nuanced sentence embeddings, improving performance across various STS tasks and demonstrating the value of incorporating backward dependencies.
IV-C2 Unsupervised Contrastive Learning Methods (UCL)
Prior to the era of LLMs, research demonstrated that utilizing data augmentation for unsupervised contrastive learning effectively enhances the quality of embeddings [158]. This unsupervised training paradigm eliminates the dependence on high-quality supervised data, improves training efficiency, and enhances scalability, thereby attracting significant attention from researchers. The emergence of LLMs, along with prompt engineering methods, can be treated as an even more efficient unsupervised learning paradigm. Nonetheless, in the era of LLMs, studies tend to employ Unsupervised Contrastive Learning Methods in embedding learning primarily as a step within multi-stage learning processes to bolster specific aspects of the final model, rather than relying solely on UCL to train the entire model as seen in previous works like SimCSE [158]. For instance, LLM2Vec [130] explores masked next-token prediction and unsupervised contrastive learning methods [158] while enabling bidirectional attention to enhance embedding performance.
IV-C3 Supervised Contrastive Learning Methods (SCL)
@@
Instruction tuning in embedding models involves adding explicit instructions to the input text to specify downstream tasks and domains or to distinguish whether the input belongs to the query or document side of a retrieval task. This allows the model to differentiate between various inputs and enhance the performance of downstream tasks. Instructor [68] first proposed the method of instruction tuning for embedding models and constructed an instruction tuning dataset containing different tasks such as classification, retrieval, and clustering. TART [53] focused on applying the instruction tuning method to multiple retrieval tasks, creating an instruction tuning training set, BERRI, which covers 40 distinct retrieval tasks. The strong instruction-following capabilities of large language models (LLMs) have further facilitated the exploration of instruction tuning in embedding models based on LLMs.  [102] explored utilizing large language models like GPT-4 to construct instruction tuning datasets and trained a decoder-only pre-trained model for text representation. Min et al. [165] introduced a curriculum learning approach for multitask instruction tuning in embeddings. InBedder [133] introduced a novel training methodology.
Furthermore, many researchers have focused on building test benchmarks to evaluate different models’ instruction-following abilities and retrieval performance, such as InstructIR [166], followIR [103], and C-STSB [167].
@@
Neither BERT nor current LLMs are primarily designed as embedding models during their pre-training phases. Consequently, there has been considerable research focused on adapting the base pre-trained language models before supervised training to make them more suitable for embedding tasks. During the BERT pre-training era, representative works include CoCondenser [168] and RetroMAE [169], which have designed distinct training objectives to enhance BERT’s text representation capabilities. In the era of LLMs, similar approaches have been extended to improve LLMs’ embedding abilities. For example, Llama2Vec [120] employs various templates that enable the model’s representations to predict which words from the original sentence and the subsequent sentence are included. This method serves as a pre-training technique to enhance the model’s capacity to capture global semantics in vector representations. Llama2Vec utilizes two auxiliary pre-training tasks to adapt the foundation model, substantially improving the model’s fine-tuned performance on a variety of dense retrieval benchmarks.
Multi-Task Learing
@@
In the supervised training phase, in addition to the contrastive learning objective, some research explores the introduction of other learning objectives to enhance the final model’s representation capabilities or to maintain the generative abilities of the LLMs, thereby enabling broader applications in downstream tasks. Piccolo2 [170] employs different loss functions for various tasks; for instance, it uses contrastive loss (CL) for retrieval tasks and margin loss for semantic textual similarity (STS) tasks. CDE [171] proposes two methods for creating contextualized document embeddings that improve neural retrieval by incorporating neighboring document information during encoding, resulting in better performance than traditional bi-encoders, particularly in out-of-domain contexts. Furthermore, GritLM simultaneously trains LLMs for both contrastive learning and Next Token Prediction (NTP) tasks, enabling the model to represent text effectively while retaining its original text generation capabilities.
IV-C4 Multi-Stage Training
@@
During the supervised learning phase, numerous studies have explored multi-stage contrastive learning training methods to enhance further the generalization and versatility of the final embedding models. For instance, influenced by earlier work from the BERT era, models like GTR [118], GTE-Qwen2 [135], and BMRetriever [122] have investigated two-stage training strategies that combine weakly supervised contrastive learning (WCL) with supervised contrastive learning (SCL). Weakly supervised contrastive learning primarily leverages a large amount of weakly supervised relevance data collected from public domains (e.g., neighboring text spans and question-answer pairs from the QA community) for training. Although this data may contain noise, the extensive domain coverage provided by typically over a billion data points significantly improves the model’s domain generalization capabilities. Additionally, NV-Retrieval proposes a phased supervised training approach tailored to different tasks (e.g., retrieval and STS tasks). In the first stage, NV-Retriever [99] utilizes only retrieval-supervised data with in-batch negatives alongside mined hard negatives. In the second stage, it blends data for retrieval tasks with datasets from other tasks (e.g., classification, regression).
IV-D Commercial Service
@@
Many AI companies provide generic text embedding services to support traditional NLP tasks and RAG. In English text embedding service, it mainly includes OpenAI  [^5] https://openai.com/index/introducing-text-and-code-embeddings, Google  [^6] https://cloud.google.com/vertex-ai/generative-ai/docs/embeddings/get-text-embeddings, Amazon  [^7] https://docs.aws.amazon.com/bedrock/latest/userguide/titan-embedding-models.html, Alibaba Cloud [^8] https://www.alibabacloud.com/help/en/model-studio/developer-reference/general-text-embedding/, Voyage  [^9] https://docs.voyageai.com/docs/embeddings and Cohere  [^10] https://cohere.com/blog/introducing-embed-v3. Due to the trade secrets involved, only a few companies have revealed some of the technology behind their services. For example, OpenAI simultaneously provides text embeddings with different dimensional settings for different levels of time-storage sensitivity in practice, where matryoshka representation learning (MRL) [172] is considered to be the key to obtaining the embedding with flexible dimensions using only one encoder. Google demonstrates that its Gecko [101], which achieves superior performance on 1B scale and 786 dimensions, relies heavily on a two-stage synthetic data generation process with LLMs: the first step generates diverse tasks and queries based on the instruction and the second step generates positive and negative samples based on the obtained tasks and queries. In addition, there have been methods [173] that attempt to distill knowledge through embeddings obtained from commercial APIs, obtaining a local embedder with similar performance successfully.
@@
TABLE V: The Overview of Text Embedding Understanding With LLMS. LM and CD represent Language Model and Context Disillation, separately. PC, ICL, RAG represent Prompt Compression, In-Context Learning and Retrieval-Augmented Generation.
@@
V Text Embedding Understanding with LLMs

### v2

@@ 4.2.2 Attention Mechanism
The discussion of attention in the current approach focuses on decoder-only LLM-based embedder. Causal Attention [200] is a common operation for decoder-based LLMs, which ensures that the language modeling can only refer to the prefix to predict the next token. However, extensive empirical studies [201, 174, 185] have shown that causal attention will degrade downstream task performance. Current approaches either (1) retain the causal attention and use other tricks (e.g., special pooling methods) or (2) convert it to bi-directional attention and make the model adapt to it.
Maintain Causal Attention
@@
If the model maintains causal attention, it is often necessary to resort to pooling tricks to mitigate the resulting negative effects. Representatives of these methods include SGPT [169], PromptEOL [166], and Echo [185]; please refer to Section 4.2.1 for details.
@@
Considering that the causal attention is implemented by a mask matrix implementation in Transformer  [^10] Transformers (https://github.com/huggingface/transformers) is the most popular open-source toolkit to instantiate LLMs., it is convenient and quick to remove the mask matrix during incremental fine-tuning. BeLLM [174] pioneered an attempt to change the causal attention of the last few layers in the decoder-only LLM to bi-directional attention. This trial is motivated by the fact that as the number of layers increases, the performance of downstream tasks does not improve monotonically, but there is a turning point. When fully converting the causal attention in decoder-only LLM to bi-directional, sufficient data or additional training tasks are required to adapt the LLM to the new attentional mechanism. The experimental results in [174] show that when training on small-scale data, changing the attention of only the last layer from causal to bi-directional can effectively improve the performance on the SemEval Benchmark; however, changing all the attentional layers directly to bi-directional will result in a significant decrease in performance. Further, NV-Embed [193] shows that LLMs can adaptively update the parameters and complete the conversion from causal attention to bi-attention without additional operations when using sufficient data (hundreds of datasets) for incremental fine-tuning. To adapt the LLM to use bi-direction attention without massive training data, LLM2Vec [164] propose a self-supervised task, masked next token prediction (MNTP). MNTP combines the ideas of mask language modeling and next token prediction, where the token is first randomly masked off from the text, and the hidden state from the previous position of the mask token is used to predict the mask token.
Dynamic Transformation
GirtLM [165] is capable of both generation and embedding tasks with a multi-task learning paradigm (Please see Section 4.3.4 for details). Therefore, GirtLM can shift to bi-directional attention for high-quality embedding while maintaining causal attention for generation, which is consistent with the training setting.
4.2.3 Additional Projector
Adding projectors is the usual strategy in text embedding. Unlike pooling strategies that operate on multiple hidden states, projectors usually operate on a single embedding for other purposes. We divide the existing work into three parts according to the purpose of the projectors.
@@ 4.2.3 Additional Projector
The hidden layer dimension (width) usually increases with the number of total parameters, according to the scaling laws [202]. The hidden states of T5-11B have 1024 dimensions, while the number is 4096 for the mainstream 7B decoder-only LLMs, which are both higher than the 768 dimensions for a bunch of BERT-based models. In practice, high-dimensional embedding dramatically increases the resource overhead for storage and inference, and a simple approach is to add a projector $g:\mathbb{R}^{d}\to\mathbb{R}^{m},m<d$ after pooling to obtain the low-dimensional embedding. With a fixed output dimension linear layer, Sentence-T5 [168] and GTR [175] empirically demonstrate that the performance on STS, classification, and retrieval tasks can be improved by scaling the embedder’s parameter only without scaling dimension. GirtLM [165] also uses a linear layer to map Mistral’s 4096-dimensional embedding to 1024 and finds that the performance of MTEB would be slightly degraded.
@@
Similar to the purpose of dimensionality reduction, converting text embedding to sparse representations can improve the inference efficiency of the model [198] and enhance its performance on partial downstream tasks such as long document retrieval [56]. For BERT-based embedders, text embeddings can be mapped to the vocabulary-length logits through the projector for mask language modeling; then, the vocabulary-size logits can be regarded as the text representation. To sparsify the representation, common methods include gate mechanism [203], top-k masking [204], and regularization terms [205]. Similarly, text embeddings from LLM-based embedders can go through the decoder layer and use a similar idea for sparsification. Specifically, PromptReps [180] uses the prompt similar to PromptEOL to obtain text embeddings and introduce ReLU function, log-saturation [198] and top-k masking to obtain the sparse representation for documents. In addition, an empirical study [206] finds that the embedding output from LLM-based embedders, which is fine-tuned by contrastive learning, still produces high-quality sparse representations via the decoding layer and simple top-$k$ masking.
@@
Search-Adaptor [207] proposes the use of additional projectors for customizing commercial text embedding interfaces, improving performance on private domain data.
4.2.4 Parameter-Efficient Fune-Tuning Module
Recall that the current parameter scales used for treating LLM-based Embedder are around 7B, e.g., both Mistral-7B and LLaMA3-8B are the most commonly used backbones. Under the popular half-precision (TF16 or BF16) setting, one 7B LLM can occupy $\sim$60GB RAM for full-parameter fine-tuning or $\sim$16G RAM for parameter-efficient fine-tuning using LoRA, which both can be accomplished on two NVIDIA V100 / A100 GPUs. Thus, whether or not parameter-efficient fine-tuning techniques are used depends primarily on the amount of data used, not on resource constraints. Some works introduce BitFit [208], or LoRA [209] as the PEFT modules and fine-tune using a single dataset, such as Wiki1M [164], SNLI [169], or MSMARCO [176], in the belief that a small-size data can stimulate the semantic generalization capabilities that LLMs themselves possess. The other works use full parameter tuning and (multiple-stage) fine-tuning based on hundreds of datasets, which was sufficient to allow large changes to the model parameters and avoid overfitting.
4.3 Optimization
@@
4.3.1 Prompt Engineering (PE)
Prompt engineering leverages the inherent capabilities of large language models without additional fine-tuning. By carefully crafting input prompts, practitioners can guide the model to produce desirable embeddings tailored to specific tasks. Several notable methods have been developed. PromptEOL [166] utilizes in-context learning to enhance sentence embedding performance. Echo [185] populates individual texts into a prompt template twice and obtains the textual embedding from the second occurrence position of the text. PromptReps [180] extends the use of prompt engineering to generate retrieval-specific embeddings. By inputting documents and queries along with relevant prompts into the LLM, it utilizes the embedding at the <EOS> position as a feature for retrieval tasks. MetaEOL [171] introduces an unsupervised text embedding technique known as Meta-Task Prompting with Explicit One-Word Limitation. In the study on PromptSTH/SUM [172], the authors introduce two novel prompt engineering techniques—Pretended Chain of Thought and Knowledge Enhancement—which enhance the expressiveness of LLM embeddings and are effective across diverse LLMs. Further, BeLLM [174] explores the impact of backward dependencies in LLMs and then proposes a backward dependency enhanced large language model (BeLLM) that transforms certain attention layers from unidirectional to bidirectional. This modification allows the model to learn more nuanced sentence embeddings, improving performance across various STS tasks and demonstrating the value of incorporating backward dependencies.
4.3.2 Unsupervised Contrastive Learning (UCL)
Before the era of LLMs, research demonstrated that utilizing data augmentation for unsupervised contrastive learning effectively enhances the quality of embeddings [76]. This unsupervised training paradigm eliminates the dependence on high-quality supervised data, improves training efficiency, and enhances scalability, thereby attracting significant attention from researchers. The emergence of LLMs, along with prompt engineering methods, can be treated as an even more efficient unsupervised learning paradigm. Nonetheless, in the era of LLMs, studies tend to employ Unsupervised Contrastive Learning Methods in embedding learning primarily as a step within multi-stage learning processes to bolster specific aspects of the final model, rather than relying solely on UCL to train the entire model as seen in previous works like SimCSE [76]. For instance, LLM2Vec [164] explores masked next-token prediction and unsupervised contrastive learning methods [76] while enabling bi-directional attention to enhance embedding performance.
4.3.3 Supervised Contrastive Learning (SCL)
@@
Instruction tuning in embedding models involves adding explicit instructions to the input text to specify downstream tasks and domains or to distinguish whether the input belongs to the query or document side of a retrieval task. This allows the model to differentiate between various inputs and enhance the performance of downstream tasks. Instructor [67] first proposed the method of instruction tuning for embedding models and constructed an instruction tuning dataset containing different tasks such as classification, retrieval, and clustering. TART [51] focused on applying the instruction tuning method to multiple retrieval tasks, creating an instruction tuning training set, BERRI, which covers 40 distinct retrieval tasks. The strong instruction-following capabilities of large language models (LLMs) have further facilitated the exploration of instruction tuning in embedding models based on LLMs.  [153] explored utilizing LLMs like GPT-4 to construct instruction tuning datasets and trained a decoder-only pre-trained model for text embedding. Min et al. [216] introduced a curriculum learning approach for multitask instruction tuning in embeddings. InBedder [191] introduced a novel training methodology. Furthermore, many researchers have focused on building test benchmarks to evaluate different models’ instruction-following abilities and retrieval performance, such as InstructIR [217], followIR [154], and C-STSB [218].
@@
Neither BERT nor current LLMs are primarily designed as embedding models during their pre-training phases. Consequently, there has been considerable research focused on adapting the base pre-trained language models before supervised training to make them more suitable for embedding tasks. During the BERT pre-training era, representative works include CoCondenser [219] and RetroMAE [220], which have designed distinct training objectives to enhance BERT’s text embedding capabilities. In the era of LLMs, similar approaches have been extended to improve LLMs’ embedding abilities. For example, Llama2Vec [176] employs various templates that enable the model’s hidden states to predict which words from the original sentence and the subsequent sentence are included. This method serves as a secondary pre-training technique to enhance the model’s capacity to capture global semantics with the hidden state from a position. Llama2Vec utilizes two auxiliary pre-training tasks, EBAE (Embedding-Based Auto-Encoding) and EBAR (Embedding-Based Auto-Regression), to adapt the foundation model, substantially improving the model’s fine-tuned performance on a variety of dense retrieval benchmarks.
4.3.4 Multi-Task Learning
@@
In the supervised training phase, in addition to the contrastive learning objective, some research explores the introduction of other learning objectives to enhance the final model’s representation capabilities or to maintain the generative abilities of the LLMs, thereby enabling broader applications in downstream tasks. Piccolo2 [221] employs different loss functions for various tasks; for instance, it uses contrastive loss (CL) for retrieval tasks and margin loss for semantic textual similarity (STS) tasks. CDE [222] proposes two methods for creating contextualized document embeddings that improve neural retrieval by incorporating neighboring document information during encoding, resulting in better performance than traditional bi-encoders, particularly in out-of-domain contexts. Furthermore, GritLM [165] simultaneously trains LLMs for both contrastive learning and Next Token Prediction (NTP) tasks, enabling the model to represent text effectively while retaining its original text generation capabilities.
4.3.5 Multi-Stage Training
@@
During the supervised learning phase, numerous studies have explored multi-stage contrastive learning training methods to enhance further the generalization and versatility of the final embedding models. For instance, influenced by earlier work from the BERT era, models like GTR [175], GTE-Qwen2 [192], and BMRetriever [178] have investigated two-stage training strategies that combine weakly supervised contrastive learning (WCL) with supervised contrastive learning (SCL). Weakly supervised contrastive learning primarily leverages a large amount of weakly supervised relevance data collected from public domains (e.g., neighboring text spans and question-answer pairs from the QA community) for training. Although this data may contain noise, the extensive domain coverage provided by typically over a billion data points significantly improves the model’s domain generalization capabilities. Additionally, NV-Retriever proposes a phased supervised training approach tailored to different tasks (e.g., retrieval and STS tasks). In the first stage, NV-Retriever [149] utilizes only retrieval-supervised data with in-batch negatives alongside mined hard negatives. In the second stage, it blends data for retrieval tasks with datasets from other tasks.
4.4 Commercial Service
@@
Many AI companies provide generic text embedding services to support traditional NLP tasks and RAG. In English text embedding service, it mainly includes OpenAI  [^11] https://openai.com/index/introducing-text-and-code-embeddings, Google  [^12] https://cloud.google.com/vertex-ai/generative-ai/docs/embeddings/get-text-embeddings, Amazon  [^13] https://docs.aws.amazon.com/bedrock/latest/userguide/titan-embedding-models.html, Alibaba Cloud [^14] https://www.alibabacloud.com/help/en/model-studio/developer-reference/general-text-embedding, Voyage  [^15] https://docs.voyageai.com/docs/embeddings and Cohere  [^16] https://cohere.com/blog/introducing-embed-v3. Due to the trade secrets involved, only a few companies have revealed some of the technology behind their services. For example, OpenAI simultaneously provides text embeddings with different dimensional settings for different levels of time-storage sensitivity in practice, where matryoshka representation learning (MRL) [223] is considered to be the key to obtaining the embedding with flexible dimensions using only one encoder. Google demonstrates that its Gecko [152], which achieves superior performance on 1B scale and 786 dimensions, relies heavily on a two-stage synthetic data generation process with LLMs: the first step generates diverse tasks and queries based on the instruction and the second step generates positive and negative samples based on the obtained tasks and queries. In addition, there have been methods [224] that attempt to distill knowledge through embeddings obtained from commercial APIs, obtaining a local embedder with similar performance successfully.
@@
Table 5: The Overview of Text Embedding Understanding With LLMS. LM and CD represent Language Model and Context Disillation, separately. PC, ICL, RAG represent Prompt Compression, In-Context Learning and Retrieval-Augmented Generation.
@@
5 Text Embedding Understanding with LLMs